OpenStudy (anonymous):

in the expression of (4r+s)^7 what is the coefficant in this term

OpenStudy (anonymous):

Which term?

OpenStudy (anonymous):

r^4 s^3

OpenStudy (anonymous):

Are you familiar with the Binomial Theorem and Pascal's Triangle ?

OpenStudy (anonymous):

a little

OpenStudy (anonymous):

I suggest a little research on it because it comes in handy. Remember that the r^4 term will also have a 4^4 term because 4 is already a coefficient of r. Basically what it tells us is that the 4^4 * r^4 * s^3 term is the fourth term (counting backwards from the exponent 7, we get : 7, 6, 5, 4 ,totaling four terms. The binomial coefficient then of the fourth term is: \[\left(\begin{matrix}7 \\ 4\end{matrix}\right) = 35\] Since we already have a 4^4 term, we multiply 4^4 * 35 = 8960 Thus the coefficient is 8960. Some research for Binomial Theorem ( http://www.purplemath.com/modules/binomial.htm)